OpenStudy (crashonce):

An AP consists of an even number of terms. The sums of the even and odd numbered terms are 300 and 240 respectively. If the first term is less than the last term by 118.5, the number of terms in the AP is ? @perl @sriramkumar

OpenStudy (crashonce):

nup doesnt really help

OpenStudy (anonymous):

Okay So The Series has even number of Terms.. And The Sum of Even terms is 300 and odd terms is 240. The even term exceeds odd term by 300-240 =60 => (n/2)d=60 => nd =120 add difference b/w the first and last terms is a+(n-1)d - a = (n-1)d = 118.5 => nd -d = 118.5 => d = 120 - 118.5 = 1.5 So you have got d; n = 120/1.5 = 80 terms.. Hope its Correct. Ref: https://answers.yahoo.com/question/index?qid=20090714051221AAY7ngO

OpenStudy (anonymous):

@CrashOnce hope this helps...